Use the pattern to identify the next two terms or figures.
3, -6, 18, -72, 360


Answer by stanbon(75887)

Use the pattern to identify the next two terms or figures.
3, -6, 18, -72, 360
------------------------------
-6 = -2*3
18 = -3*-6
-72= -4*18
360 = -5*72
------
Then -6*360 = -2160
And -7*-2160 = 15120
==========================
Cheers,
Stan H.
=============